# Adapting Conformal Prediction to Distribution Shifts Without Labels

### v1

@@ 5.1 Stationary shifts
Figure 4: Our $\texttt{E{\small A}CP}_{2}$ method is able to improve coverage using various neural network models and architectures, as well as a large and diverse range of distribution shifts. It consistently “hugs” the desired coverage rate, while maintaining practical set sizes. Results are averaged across five severity levels for each corruption type in the ImageNet-C dataset. We zoom in on the right to clearly see the benefit of adapting at test time on coverage and set sizes.
We report performance using a number of neural networks architectures and sizes on all 19 corruption types averaged across severities for each corruption in Figure 4. Here, the methodology and hyper-parameters are exactly the same for all models, including the vision transformers. We can see that even with no hyper-parameter tuning (including on the TTA algorithm), our method considerably reduces the coverage gap across a variety of neural networks. An immediate observation is that while ETA does not recover coverage, it also consistently leads to smaller set sizes. While minimizing set sizes is indeed desirable, this may be of secondary importance to the primary goal of achieving the desired coverage level. Further, we note that the set sizes without ECP are in fact consistent and with little variance, regardless of the empirical coverage. We argue that this is actually an undesirable quality. As an uncertainty quantification framework, conformal prediction is considered to convey uncertainty through its set sizes; greater uncertainty is supposed to be reflected in larger sets, and vice versa. Our results suggest that explicitly incorporating uncertainty into the conformal predictor seems necessary to accurately convey the underlying uncertainty under distribution shift. While even large mis-coverage still results in small set sizes using ETA, larger sets under EACP consistently correspond to lower coverage — an important and practical quality for end-users.
5.2 Continuous shifts
@@
Table 3: We evaluate performance on the challenging setting of continuously shifting distributions. The “label free” column denotes whether a method relies on labels at test-time from the target data. We recall that SplitCP does not adapt to new data. In addition to the average coverage ($\uparrow$) and average size ($\downarrow$), we also measure the worst local corruption error LCE128 ($\downarrow$) and worst local set size LSS128, ($\downarrow$) on a sliding window of 128 test points.
@@
SAOCP (Seedat et al., 2023) 0.79 145 0.24 353 0.78 139 0.28 349
FACI (Gibbs and Candes, 2021) 0.90 101 0.07 455 0.90 142 0.09 450
@@
We further investigate the setting of continuous distribution shift in Table 3. This setting is similar to that studied in previous Online Conformal Prediction (OCP) works (Zhang et al., 2024; Bhatnagar et al., 2023b; Gibbs and Candès, 2023; Gibbs and Candes, 2021). This involves shifting between ImageNet-C severity level 1 to level 5 (either suddenly or gradually), while sampling random corruptions at each corresponding severity. We emphasize that this is a particularly challenging setting, as it presents a continuous shift in both the magnitude as well as type of corruption. See Appendix A.2 for more details on this set-up. In addition to unsupervised methods, we also compare with supervised OCP methods that rely on the correct label being revealed after every prediction. We notice that on gradual shifts in severity, $\texttt{E{\small A}CP}_{2}$ achieves nearly the same coverage results as these strong baselines, with greatly reduced set-sizes. We believe this considerable reduction in set size is due to two factors. First, $\texttt{E{\small A}CP}_{2}$ is able to update the underlying neural network on new data. Second, while OCP is able to learn the (potentially expansive) necessary threshold, $\texttt{E{\small A}CP}_{2}$ instead directly updates the conformal scores themselves. In addition to average coverage and set size, we also measure the local coverage error LCE128 across the worst sliding window of 128 samples, and similarly the worst local set size, LSS128. While the supervised methods unsurprisingly result in better local coverage, they also lead to local set sizes that are an order of magnitude larger than those produced by $\texttt{E{\small A}CP}_{2}$.
@@ 6 Conclusion
This paper studies how to adapt CP methods to distribution-shifted data, without relying on labels from the target dataset. This is an important challenge in many real world settings, where exchangeability assumptions are violated and labels may be difficult to attain. We propose an uncertainty-aware method based on the prediction entropy (ECP), and leverage unsupervised test time adaptation to update the base model and refine its uncertainty (EACP). We demonstrate that the proposed methods are able to recover the desired error rate on a wide range of distribution shifts, while maintaining efficient set sizes. Furthermore, they are even competitive with supervised approaches on challenging and continuously shifting distributions. We hope this inspires future works continuing to tackle this important challenge.
Appendix
Appendix A contains detailed information on our studied datasets and experimental protocols, including TTA hyper-parameters, CP procedure, and continuous shift set-up. Appendix B discusses other possible uncertainty measures and their deficiencies. Appendix C presents results with another base TTA method, Appendix D has results of more neural network architectures on natural distribution shifts, Appendix E has graphs performance on continuous-shifts at each time-step, and Appendix F has full ImageNet-C results on each severity level.
Appendix A Experimental details
A.1 Dataset details
@@
  • ImageNet-V2 [Recht et al., 2019] is a new ImageNet test-set that contains 10,000 images that were collected by closely following the original ImageNet data collection process.
@@ A.2 Experimental protocols
Our split conformal prediction set-up follow previous works [Angelopoulos et al., 2021, Angelopoulos and Bates, 2023], which divides a held-out dataset into a calibration and test set. On ImageNet variants, we split the original validation set in half to produce 25,000 calibration points and 25,000 in-distribution test points. The calibrated scores and / or threshold are then used for subsequent distribution-shifted data. On the WILDS datasets, we similarly split the in-distribution validation sets.
@@
We adopt a slightly modified version of the experimental design for continuous distribution shift presented in previous works [Bhatnagar et al., 2023a, Zhang et al., 2024]. This involves sampling random corruptions from the ImageNet-C dataset under two regimes: gradual shifts where the severity level first increases in order from $\{1,...,5\}$ then decreases from $\{5,...,1\}$, and sudden shifts where the severity level alternates between 1 and 5. In addition to sampling random corruptions, we also consider in Figure 7 results on the “easier” setting of shifting severity’s on a single corruption type.
Appendix B What is the right measure of uncertainty?
@@
Appendix C Other TTA methods
@@
Appendix D More architecture comparisons
@@ Appendix E Continuous shifts
Appendix F ImageNet-C all severity levels
@@
Figure 8: Performance on 19 ImageNet-C corruptions on each severity level. $\texttt{E{\small A}CP}_{2}$ hugs the desired coverage line on nearly all severity levels.

### v2

@@ 4 Experiments
Figure 4: Our $\texttt{E{\small A}CP}_{2}$ method is able to improve coverage using various neural network models and architectures, under a diverse range of distribution shifts. It consistently “hugs” the desired coverage rate, while maintaining practical set sizes. Results are averaged across five severity levels for each corruption type in the ImageNet-C dataset. We zoom in on the right to clearly show the benefit of adapting at test time. Larger markers reflect a larger neural network parameter count.
@@
Table 3: We evaluate performance on the challenging setting of continuously shifting distributions. The “label free” column denotes whether a method relies on labels at test-time from the target data. We recall that SplitCP does not adapt to new data. In addition to the average coverage ($\uparrow$) and average size ($\downarrow$), we also measure the worst local coverage error LCE128 ($\downarrow$) and worst local set size LSS128, ($\downarrow$) on a sliding window of 128 test points.
@@
SAOCP (Bhatnagar et al., 2023a) 0.79 145 0.24 353 0.78 139 0.28 349
DtACI (Gibbs and Candès, 2024) 0.90 101 0.07 455 0.90 142 0.09 450
@@
Next, Figure 4 contains the results using neural networks of various architectures and parameter counts, on all 19 corruption types of ImageNet-C (average across five severity levels). Besides showing the superior performance of our methods, we observe that the SplitCP baseline (with and without TTA) generates prediction sets with little variance in the set sizes, regardless of the achieved coverage rates. We argue that this is an undesirable behavior, as the set sizes themselves are often used to encode uncertainty evaluations by set-valued classifiers. Our results demonstrate that explicitly incorporating the base model’s own uncertainty into CP can help mitigating this issue.
Continuous shifts
Finally, we investigate continuous distribution shifts, and the results are shown in Table 3. This has been previously studied under online conformal prediction, and we build on the experimental setup of (Bhatnagar et al., 2023b; Zhang et al., 2024). Specifically, the environment shifts between ImageNet-C severity level 1 to level 5 (either suddenly or gradually; see Figure 7), while sampling random corruptions at each corresponding severity. We emphasize that this is a particularly challenging task, as it presents a continuous shift in both the magnitude as well as type of corruption. See Appendix B.2 for more details on this experiment. Here, we compare with existing supervised methods that rely on the correct label being revealed after every prediction.
Overall, our methods demonstrate competitive performance with respect to supervised baselines: the average set sizes are significantly smaller despite a slight drop in the average coverage rate. We also measure the local coverage error LCE128 across the worst sliding window of 128 samples, and similarly the worst local set size, LSS128. While the supervised methods unsurprisingly result in better local coverage, they also lead to local set sizes that are much larger.
@@ 5 Conclusion
This paper studies how to improve set-valued classification methods on distribution-shifted data, without relying on labels from the target dataset. This is an important challenge in many real world settings, where exchangeability assumptions are violated and labels may be difficult to attain. We propose an uncertainty-aware method based on the prediction entropy (ECP), and leverage unsupervised test time adaptation to update the base model and refine its uncertainty (EACP). We demonstrate that the proposed methods are able to recover the desired error rate on a wide range of distribution shifts, while maintaining efficient set sizes. Furthermore, they are even competitive with supervised approaches on challenging and continuously shifting distributions. We hope this inspires future works continuing to tackle this important challenge.
Acknowledgements.
We appreciate the feedback from anonymous reviewers. ZZ and HY are partially funded by Harvard University Dean’s Competitive Fund for Promising Scholarship. KK and GWT acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) and the Ontario Ministry of Economic Development and Innovation under the Ontario Research Fund–Research Excellence (ORF-RE) program. Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute.
Appendix A Additional related works
CP in decision making
Our interest in the considered setting – distribution shifts without test time labels – is mainly motivated by the growing applications of CP in autonomous decision making. A very much incomplete list: see [Lekeufack et al., 2024] for a generic treatment; [Lindemann et al., 2023] for trajectory optimization in robotics; [Yang and Pavone, 2023, Gao et al., 2024] for 3D vision; [Kumar et al., 2023, Cherian et al., 2024, Gui et al., 2024, Mohri and Hashimoto, 2024, Quach et al., 2024] for large language models (LLMs); and [Ren et al., 2023] for LLM-powered robotics.
CP under distribution shifts
As discussed in the main paper, considerable efforts have been devoted to developing CP methods robust to distribution shifts. We now survey two possible directions and their respective limitations.
  • The first direction does not require test time labels, but the distribution shift is assumed to be simple in some sense. For example, Tibshirani et al. [2019] studied CP under covariate shifts, where the distribution of the label $y$ conditioned on the covariate $x$ remains unchanged. Here, it suffices to use the classical likelihood ratio reweighting on the calibration dataset, but accurately estimating the likelihood ratio can be challenging in practice. Another idea is to take a robust optimization perspective by assuming a certain maximum level of distribution shift and protecting against the worst case, e.g., [Roth, 2022, Chapter 8] and [Cauchois et al., 2024]. The weakness here is the sensitivity to the hyperparameter, and the obtained prediction sets could be overly conservative.
    Various works built on these two ideas. Barber et al. [2023] generalized the reweighting idea to handle mild but general distribution shifts, but choosing the weights is generally unclear in practice. Ai and Ren [2024] tackled general distribution shifts by combining reweighting and robust optimization, which also combines the strengths and limitations from the two sides. Ge et al. [2024] extended the two ideas to the aggregation of multiple CP algorithms.
  • The second direction is connecting CP to adversarial online learning. A line of works [Gibbs and Candes, 2021, Angelopoulos et al., 2023, Gibbs and Candès, 2024, Bhatnagar et al., 2023b, Zhang et al., 2024] applied regret minimization algorithms in OCO to select the score threshold in CP, and Bastani et al. [2022] achieved this task using multicalibration. By relaxing the CP objective from the coverage probability to the post-hoc coverage frequency, these methods can handle arbitrary continual distribution shifts. However, they require the true label to be provided after every prediction, which is a limiting requirement for many use cases in autonomous decision making. Our experiments will show that it is possible to achieve comparable performance in these settings without this limitation, i.e., being “label free”.
Appendix B Experimental details
B.1 Dataset details
@@
  • ImageNet-V2 [Recht et al., 2019] is an ImageNet test-set that contains 10,000 images that were collected by closely following the original ImageNet data collection process.
@@ B.2 Experimental protocols
Our split conformal prediction set-up follows previous works [Angelopoulos et al., 2021, Angelopoulos and Bates, 2023], which divides a held-out dataset into a calibration and test set. On ImageNet variants, we split the original validation set in half to produce 25,000 calibration points and 25,000 in-distribution test points. The calibrated scores and / or threshold are then used for subsequent distribution-shifted data. On the WILDS datasets, we similarly split the in-distribution validation sets.
@@
We adopt a slightly modified version of the experimental design for continuous distribution shift presented in previous works [Bhatnagar et al., 2023a, Zhang et al., 2024]. This involves sampling random corruptions from the ImageNet-C dataset under two regimes: gradual shifts where the severity level first increases in order from $\{1,...,5\}$ then decreases from $\{5,...,1\}$, and sudden shifts where the severity level alternates between 1 and 5. In addition to sampling random corruptions, we also consider in Figure 7 results on the “easier” setting of shifting severities on a single corruption type.
Appendix C What is the right measure of uncertainty?
@@
Appendix D Additional experiments
D.1 Other TTA methods
@@
D.2 More architecture comparisons
@@ D.3 Continuous shifts
D.4 ImageNet-C all severity levels
@@
Figure 8: Performance on 19 ImageNet-C corruptions on each severity level. $\texttt{E{\small A}CP}_{2}$ hugs the desired coverage line on nearly all severity levels. Larger markers indicate larger parameter count.
D.5 Oracle results
Here we compare our methods with an oracle that has observed labels from the distribution-shifted dataset. Specifically, the oracle is the THR conformal prediction method [Sadinle et al., 2019] that has been calibrated on half of the distribution-shifted dataset, following regular split conformal. Since the oracle is guaranteed to provide the desired coverage level in this set-up, our comparison focuses on the prediction set sizes; we refer to the main paper for coverage comparisons. We observe in Table 6 that in every case except FMOW, a variant of ECP and EACP achieves smaller set sizes than the oracle. In Table 7, EACP consistently achieves substantially smaller set sizes on ImageNet-C while also recovering error targets (see Table 2). We reiterate here that smaller sets are preferred if error rates are maintained.
Table 6: ECP and EACP achieve prediction set sizes that are often equal or smaller than the oracle method. Coverage rate is 0.90.
Method ImageNet-V2 ImageNet-R ImageNet-A iWildCam RXRX1 FMOW
Set Size ORACLE 6.8 79.0 95.3 6.6 140 7.87
$\texttt{ECP}_{1}$ 4.2 9.1 7.4 3.8 105 10.3
$\texttt{ECP}_{2}$ 7.6 23.3 15.1 5.5 137 15.3
$\texttt{E{\small A}CP}_{1}$ 4.5 6.8 8.7 3.7 133 11.1
$\texttt{E{\small A}CP}_{2}$ 8.7 16.1 10.1 5.6 177 16.4
Table 7: Comparison of ECP and EACP on a subset of synthetic shifts. The numbers refer to severity level.
Method Contrast Brightness Gaussian Noise Motion Blur
1 3 5 1 3 5 1 3 5 1 3 5
ORACLE 5.5 30.3 562 2.5 3.7 9.8 6.2 70.6 317 9.7 101 638
Set Size $\texttt{ECP}_{2}$ 10.5 27.8 180 5.3 7.7 14.9 10.0 43.1 109 12.9 43.7 79.0
$\texttt{E{\small A}CP}_{2}$ 5.5 7.4 25 4.5 5.7 7.6 5.7 16.0 42.7 6.3 12.8 25.5
D.6 In-distribution results
In Table 8, we observe that our methods maintain coverage and reasonable set sizes on in-distribution data.
Table 8: Results on in-distribution data using ImageNet-1k validation set.
SplitCP $\texttt{ECP}_{1}$ $\texttt{ECP}_{2}$ $\texttt{E{\small A}CP}_{1}$ $\texttt{E{\small A}CP}_{2}$
Coverage 0.90 0.92 0.94 0.91 0.93
Set size 2.1 2.8 4.2 2.8 4.1
D.7 Comparison with weighted CP
Tibshirani et al. [2019] present a method for improving coverage under covariate shift by re-weighing calibration scores based on an estimated likelihood ratio (wcp). Although estimating likelihood ratios in our setting is challenging, we nonetheless present a comparison here for completeness. We follow their approach and train a probabilistic classifier, here a CNN, on each calibration-test pair.
Table 9 suggests that this method may have limited performance in our studied setting. This may be due to the challenge in estimating accurate likelihood ratios in high-dimensional settings, [Cauchois et al., 2024]. We do not claim that wcp definitely cannot perform well, however the sparsity of previous literature here suggests that further studies may be required. Finally, note that wcp is ill-suited for the case of continuously shifting distributions, further limiting its general applicability.
Table 9: The wcp method appears to provide minimal coverage improvements in this setting, possibly due to the difficulty in estimating likelihood ratios.
Method ImageNet-V2 ImageNet-R ImageNet-A
Coverage SplitCP 0.81 0.50 0.03
wcp 0.82 0.35 0.06
$\texttt{ECP}_{2}$ 0.91 0.72 0.27
$\texttt{E{\small A}CP}_{2}$ 0.91 0.80 0.30
Set Size SplitCP 2.5 3.4 3.4
wcp 2.6 0.74 4.3
$\texttt{ECP}_{2}$ 7.6 23.3 15.1
$\texttt{E{\small A}CP}_{2}$ 8.7 16.1 10.1
D.8 Affects of Model Calibration
We conduct experiments investigating affects of model calibration on the robustness of our entropy-based method. We employ temperature scaling with $T<1.0$ to deliberately sharpen the model’s logits, inducing overconfidence and emulating a more poorly calibrated base model. We further compare with a temperature value determined to improve calibration (on in-distribution data), as measured using Expected Calibration Error (ECE).
Results for two distribution-shifted datasets can be seen in Table 10(b). Our results suggest that our method is fairly robust to miscalibration. Considerably higher ECE (worse calibration) leads to only minor drops in prediction set accuracy. This may potentially be due to our use of an entropy quantile, which aggregates uncertainty across the test data, mitigating issues caused by a smaller number of miscalibrated points.
Table 10: Our use of an entropy quantile renders our method robust to poor model calibration.
(a) ImageNet-V2 coverage results with different temperature scaling.
Method $T=1.15$ $T=1.0$ $T=0.9$ $T=0.8$
ECE = 0.02 ECE = 0.03 ECE = 0.06 ECE = 0.08
SplitCP 0.81 0.81 0.81 0.81
ETA 0.81 0.81 0.81 0.81
ECP 0.92 0.91 0.90 0.87
EACP 0.92 0.91 0.90 0.88
(b) ImageNet-R coverage results with different temperature scaling.
Method $T=1.15$ $T=1.0$ $T=0.9$ $T=0.8$
ECE = 0.02 ECE = 0.03 ECE = 0.06 ECE = 0.08
SplitCP 0.49 0.50 0.50 0.51
ETA 0.61 0.62 0.62 0.63
ECP 0.76 0.72 0.70 0.67
EACP 0.83 0.80 0.77 0.75